5km + 30m + 7cm=
1
Expert's answer
2019-03-29T11:04:32-0400

this is 5km + 30m + 7cm = 5,03007 km

or 5km + 30m + 7cm = 5030,07 m

or 5km + 30m + 7cm = 503007 cm